# Large-$N$ SU(4) Schwinger boson theory for coupled-dimer antiferromagnets

### v2

@@ Abstract
We develop a systematic large-$N$ expansion based on the Schwinger boson representation of SU(4) coherent states of dimers for the paradigmatic spin-$1/2$ bilayer square lattice Heisenberg antiferromagnet. This system exhibits a quantum phase transition between a quantum paramagnetic state and a Néel order state, driven by the coupling constant $g=J^{\prime}/J$, which is defined as the ratio between the inter-dimer $J^{\prime}$ and intra-dimer $J$ exchange interactions. We demonstrate that this approach accurately describes static and dynamic properties on both sides of the quantum phase transition. The critical coupling constant $g_{c}\approx 0.42$ and the dynamic spin structure factor reproduce quantum Monte Carlo results with high precision. Notably, the $1/N$ corrections reveal the longitudinal mode of the magnetically ordered phase along with the overdamping caused by its decay into the two-magnon continuum. The present large-$N$ $SU(N)$ Schwinger boson theory can be extended to more general cases of quantum paramagnets that undergo a quantum phase transition into magnetically ordered states.
@@ I Introduction
Several other approaches have been employed to solve the bilayer antiferromagnet  [17, 18, 19, 20, 21, 22, 23, 24, 25, 16, 26, 27, 28]. From the magnetically ordered side, conventional linear spin-wave theory and its modified self-consistent versions, along with the Schwinger boson (SB) $SU(2)$ mean-field approximation, have incorrectly predicted a weak first-order transition at much smaller critical values of $g_{c}$ (see Table  1), in stark contrast with QMC results. Chubukov and Morr [17] identified that the failure of these theories was due to the neglect of longitudinal fluctuations, specifically those along the local magnetization, which are expected to be significantly enhanced near the QPT.
@@
The four bond operators can be identified with the four Schwinger bosons associated with the fundamental irreducible representation (irrep) of SU(4) [30], as their bilinears faithfully represent the $\mathrm{SU}(4)$ generators of infinitesimal unitary transformations within the local four-dimensional Hilbert space of each dimer. This representation can be extended to any completely symmetric irrep of $\mathrm{SU}(4)$ by adjusting the local constraint from one to $M$ bosons per dimer [35], where $M$ is a positive integer that labels different irreps. Notably, the classical limit of this theory is obtained by taking the large $M$ limit [36]. In this context, $M$ plays a role analogous to the spin value $S$ in the conventional SU(2) case.
@@
The SU(2) Schwinger boson theory can be improved by applying a Gutzwiller projection [38] to enforce the local constraint. This procedure results in a better critical value of $g_{c}$ than the mean-field result, although $g_{c}$ remains significantly smaller than the QMC result (see Table 1). Another possibility is to perform a $1/N$-correction of the mean-field (saddle-point) solution. However, both, the Gutzwiller projection and large-$N$ expansion, start from an SU(2) Schwinger boson representation of the spin $S=1/2$ operators.
@@
Table 1: Representative numerical and analytical results for the bilayer Heisenberg quantum antiferromagnet. $g_{c}$ indicates the critical coupling, “Transition” the order of the phase transition, "Goldstone“ indicates if the theory captures or not the Goldstone modes of the magnetically ordered phase. QMC stands for Quantum Monte Carlo, LSWT for linear spin wave theory, MSWT for modified spin wave theory, SBMFT for SU(2) Schwinger boson mean-field theory, SB Gutzwiller for SBMFT Gutzwiller projected, BOT-$\lambda$ for bond operator mean-field theory with the constraint imposed by a Lagrange multiplier, BOT+Brueckner for bond-operator theory with a constraint of no triplet double-occupancy, large-$d$ BOT for the $1/d$ expansion of the BOT, SE for dimer series expansions, CST for continuous similarity transformations, large-$M$ SU(4) HP for the generalized Holstein-Primakoff spin wave theory (equivalent to large-$d$ BOT for $d=\infty$), and large-$N$ SU(4) SB corresponds to this work.
@@
In this work, we develop a systematic large-$N$ expansion based on the $\mathrm{SU}(4)$ coherent state description of a single dimer. This approach generalizes the original dimer problem with two antiferromagnetically coupled $\mathrm{SU}(2)$ spins on each site of the dimer to two antiferromagnetically coupled $\mathrm{SU}(n)$ spins, with $n\geq 2$. Unlike the traditional SB approximation, which usually employs $\mathrm{SU}(n)$ Schwinger bosons to represent the $\mathrm{SU}(n)$ spin components [41], we start from $\mathrm{SU}(N=n^{2})$ SBs to include all quantum mechanical states of the generalized dimer in the manifold of $\mathrm{SU}(N)$ coherent states. Our theory gives rise to a controlled structure and a $1/N$ expansion of the BOT mean-field approximation, where the boson number constraint is imposed by a Lagrange multiplier.
@@ II $\mathrm{SU}(4)$ coherent-state description of a dimer
where $S$ labels the irrep of $\mathrm{SU}(2)$ that determines the matrix form of the generators given in Eq. (3). Note that $S=1/2$ for the particular case of interest. However, as we mentioned in the Introduction, a large-$N$ approximation based on these SBs is not quantitatively accurate because the path-integral formulation is parametrized in terms of direct products of $\mathrm{SU}(2)$ coherent states. In other words, the intra-dimer entanglement must be built in via quantum fluctuations that favor linear combination of these product states.
@@
provide a faithful representation of generators of $\mathrm{SU}(4)$ in the completely symmetric irreps labelled by the integer $M$ (e.g. $M=1$ for the fundamental irrep of $\mathrm{SU}(4)$). Note that unlike the $\mathrm{SU}(2)$ SBs, these bosons have only a dimer index $j$ because they create all the possible quantum mechanical states of each dimer. In other words, the $\mathrm{SU}(4)$ SBs create coherent states of the $\mathrm{SU}(4)$ Lie algebra (for completely symmetric irreps) that span the CP3 manifold of quantum mechanical states of a four-level quantum mechanical system. To make contact with our problem of interest, we can regard the single dimer $j$ as the four-level system, set $M=1$ (fundamental irrep of $\mathrm{SU}(4)$) and choose the basis of $\mathrm{SU}(4)$ SBs, such that the $\mu=0$ boson creates the singlet state:
@@
The important difference between the $\mathrm{SU}(2)$ and the $\mathrm{SU}(4)$ SBs is that coherent states of the latter include intra-dimer singlet states, meaning that the intra-dimer entanglement is already built-in the $\mathrm{SU}(4)$ coherent states of the path-integral formulation and only the inter-dimer entanglement must be generated via the inclusion of quantum fluctuations. As expected from this simple observation, we will see in the next sections that the $\mathrm{SU}(4)$ SB theory (large-$N$ expansion) is indeed much more accurate than the $\mathrm{SU}(2)$ SB theory for low-dimensional coupled-dimer systems.
@@ III $\mathrm{SU}(4)$ Schwinger boson theory
In terms of the $\mathrm{SU}(4)$ SB, the single-dimer Hamiltonian takes the diagonal form:
@@
Given the local constraint in Eq. (5), the ground state $|\Psi_{0}\rangle$ of ^$\hat{\mathscr{H}}_{0}$ (direct product of singlet states) satisfies $\hat{b}_{j,0}^{\dagger}\hat{b}_{j,0}|\Psi_{0}\rangle=|\Psi_{0}\rangle$ for $\forall j$.
The inter-dimer interaction $\hat{{\mathscr{H}}}^{\prime}$ generates quantum fluctuations that convert a pair of singlet states into pairs of triplets that form a inter-dimer singlet. In addition, the low-energy excitations that were localized triplet modes in absence of $\hat{{\mathscr{H}}}^{\prime}$, become extended triplon modes because finite inter-dimer coupling can exchange a singlet and a triplet state in two different dimers. Both effects, creation of virtual triplon pairs and effective triplon hopping produce a finite dispersion of the still gapped single-triplon mode.
@@
$\hat{\mathscr{H}}^{\prime}$ (14) is expressed in a form that is manifestly SU(2) symmetric. Physically, the different terms of $\hat{\mathscr{H}}^{\prime}$ can be interpreted in the following way: $\hat{A}^{\dagger}\hat{S}$ represents the process of destroying two neighboring singlets followed by the creation of two triplets, which together form a singlet (${\rm h.c.}$ corresponds to the inverse process); $\hat{T}^{\dagger}\hat{B}$ the exchange between singlet and triplet neighbours, $:\hat{B}^{\dagger}\hat{B}:$ the exchange between neighboring triplets, and $\hat{A}^{\dagger}\hat{A}$ projects onto singlet formed between two neighboring triplets.
@@ III.1 Large-$N$ generalization
To study this interacting boson system, we use the large-$N$ technique, which extends the original $\mathrm{SU}(2)$ invariant system to a system with $\mathrm{SU}(n)$ symmetry. Each dimer consists of two antiferromagnetically coupled $\mathrm{SU}(n)$ spins described by:
^$\hat{\mathscr{H}}_{0}=\frac{4\tilde{J}(N)}{N}\sum_{j}\left(\hat{O}_{j,+}^{\mu}\hat{\tilde{O}}_{j,-}^{\mu}+C(N)\right),$ (22)
where $N=n^{2}$, $\hat{O}^{\mu}_{j,+}$ denotes the generators of the $\mathrm{SU}(n)$ group at the “$+$” site of the dimer in the fundamental irrep, while $\hat{\tilde{O}}_{j,-}^{\mu}$ are the $\mathrm{SU}(n)$ generators in the conjugate representation for the “--” site. $C(N)$ is an irrelevant constant that is used to set the highest energy eigenvalue to zero. The tensor product of these representations is reducible, i.e., ${n}\otimes\bar{{n}}={1}\oplus{n^{2}-1}$, where the numbers refer to the dimensions of the corresponding irrep. We associate the $N$ flavors of the bosons with the energy levels of each dimer, denoted as $\hat{b}_{j,\sigma}$, where $\sigma=0,1,2,...,N-1$. Here, $\hat{b}_{j,0}^{\dagger}$ creates an $\mathrm{SU}(n)$ singlet state with energy $-J$ that scales as $N^{0}$ in the large-$N$ limit. The remaining bosons generate an $N-1$ dimensional multiplet with zero energy, assuming appropriate choices of the generalized exchange constant $\tilde{J}(N)$ and the global shift $C(N)$. The $\mathrm{SU}(n)$ generators are given by:
@@
where $11\leq\mu\leq N-$ 1 $,f_{\mu\nu\rho}$ are the structure constants of the $\mathfrak{su}(n)$ Lie algebra (completely antisymmetric under exchange of any pair of indices), and $d_{\mu\nu\rho}$ is a fully symmetric third rank tensor, defined by the anticommutator of the $\mathrm{SU}(n)$ generators:
$d_{\mu\nu\rho}=2\operatorname{Tr}\left[\left\{T^{\mu},T^{\nu}\right\}T^{\rho}\right],$ (25)
where the index $11\leq\mu,\nu,\rho\leq N-$ 1. For the $\mathrm{SU}(2)$ case, $d_{\mu\nu\rho}\equiv 0$ and $f_{\mu\nu\rho}=\epsilon_{\mu\nu\rho}$, so we recover the expression for spin operators in Eq. (12).
In this generalization of the model to arbitrary $n$ values, with $N=n^{2}$, the isolated dimer Hamiltonian retains the form of the original spin-$1/2$ Hamiltonian, as defined in Eq. (13). The extension of the inter-dimer interaction is based on Eq. (14), namely,
$+\frac{2J^{\prime}}{N}\sum_{j,\delta}\left(:\hat{B}_{j,\delta}^{\dagger}\hat{B}_{j,\delta}:-\hat{A}_{j,\delta}^{\dagger}\hat{A}_{j,\delta}\right),$
where the link operators
$\hat{S}_{j,\delta}=\hat{b}_{j,0}\hat{b}_{j+\delta,0},\quad\hat{T}_{j,\delta}=\hat{b}_{j,0}^{\dagger}\hat{b}_{j+\delta,0},$ (27)
are bilinears in singlet bosons, while the link operators
$\hat{A}_{j,\delta}=\sum_{\mu=1}^{N-1}\hat{b}_{j,\mu}\hat{b}_{j+\delta,\mu},\quad\hat{B}_{j,\delta}=\sum_{\mu=1}^{N-1}\hat{b}_{j,\mu}^{\dagger}\hat{b}_{j+\delta,\mu},$ (28)
consist of bilinears in the $N-1$ flavor bosons that transform according to the (self-conjugate) adjoint irrep of $\mathrm{SU}(N)$. Notably, all four bond operators are $\mathrm{SU}(n)$ invariant.
Lastly, the local constraint Eq. (5) is generalized to:
$\sum_{\mu=0}^{N-1}\hat{b}^{\dagger}_{j,\mu}\hat{b}_{j,\mu}=M,$ (29)
where the integer $M$ labels different completely symmetric irreps of the $\mathrm{SU}(N)$ group. The $M\to\infty$ limit corresponds to a classical limit based on $\mathrm{SU}(N)$ coherent states of the extended model and quantum corrections can be incorporated via a $1/M$ expansion [36, 43, 44]. In the large-$N$ approach discussed here, $M$ must be proportional to $N$ and the ratio $\kappa=M/N$ remains a finite number in the large-$N$ limit (see Fig. 2). In this large-$N$ limit, the average boson density for each flavor remains of order one. For our original problem, we have $N=4$ and $M=1$.

### v3

@@ Abstract
We develop a systematic large-$N$ expansion based on the Schwinger boson representation of SU(4) coherent states of dimers for the paradigmatic spin-$1/2$ bilayer square lattice Heisenberg antiferromagnet. This system exhibits a quantum phase transition between a quantum paramagnetic state and a Néel order state, driven by the coupling constant $g=J^{\prime}/J$, which is defined as the ratio between the inter-dimer $J^{\prime}$ and intra-dimer $J$ exchange interactions. We demonstrate that this approach accurately describes static and dynamic properties on both sides of the quantum phase transition. The critical coupling constant $g_{c}\approx 0.42$ and the dynamic spin structure factor reproduce quantum Monte Carlo results with high precision. Notably, the $1/N$ corrections reveal the longitudinal mode of the magnetically ordered phase along with the overdamping caused by its decay into the two-magnon continuum. The present large-$N$ SU(4) Schwinger boson theory can be extended to more general cases of quantum paramagnets that undergo a quantum phase transition into magnetically ordered states.
@@ I Introduction
Several other approaches have been employed to solve the bilayer antiferromagnet  [17, 18, 19, 20, 21, 22, 23, 24, 25, 16, 26, 27, 28]. From the magnetically ordered side, conventional linear spin-wave theory and its modified self-consistent versions, along with the Schwinger boson (SB) SU(2) mean-field approximation –that is, the large-$N$ SU(2) SB theory–, have incorrectly predicted a weak first-order transition at much smaller critical values of $g_{c}$ (see Table  1), in stark contrast with QMC results. Chubukov and Morr [17] identified that the failure of these theories was due to the neglect of longitudinal fluctuations, specifically those along the local magnetization, which are expected to be significantly enhanced near the QPT.
@@
The four bond operators can be identified with the four Schwinger bosons associated with the fundamental irreducible representation (irrep) of SU(4) [30, 35, 36], as their bilinears faithfully represent the $\mathrm{SU}(4)$ generators of infinitesimal unitary transformations within the local four-dimensional Hilbert space of each dimer. This representation can be extended to any completely symmetric irrep of $\mathrm{SU}(4)$ by adjusting the local constraint from one to $M$ bosons per dimer [37], where $M$ is a positive integer that labels different irreps. Notably, the classical limit of this theory is obtained by taking the large $M$ limit [38]. In this context, $M$ plays a role analogous to the spin value $S$ in the conventional SU(2) case.
@@
The SU(2) Schwinger boson theory can be refined by applying a Gutzwiller projection [40] to enforce the local constraint. This approach yields a more accurate critical value of $g_{c}$ compared to the mean-field result, although remains significantly lower than the QMC estimate (see Table 1). A more sophisticated variational scheme [41] incorporating the Gutzwiller projection can further improve the accuracy of $g_{c}$; however, it fails to correctly capture the gapless Goldstone modes in the AFM phase. An alternative approach is to include $1/N$ corrections. However, both the Gutzwiller projection and the large-$N$ expansion originate from an SU(2) Schwinger boson representation of the spin-$S=1/2$ operators.
@@
Table 1: Representative numerical and analytical results for the bilayer Heisenberg quantum antiferromagnet. $g_{c}$ indicates the critical coupling, “Transition” the order of the phase transition, "Goldstone“ indicates if the theory captures or not the Goldstone modes of the magnetically ordered phase. QMC stands for Quantum Monte Carlo, LSWT for linear spin wave theory, MSWT for modified spin wave theory, large-$N$ SU(2) SB for the SU(2) Schwinger boson mean-field theory, SB Gutzwiller for SBMFT Gutzwiller projected, BOT-$\lambda$ for bond operator mean-field theory with the constraint imposed by a Lagrange multiplier, BOT+Brueckner for bond-operator theory with a constraint of no triplet double-occupancy, large-$d$ BOT for the $1/d$ expansion of the BOT, SE for dimer series expansions, CST for continuous similarity transformations, large-$M$ SU(4) HP for the generalized Holstein-Primakoff spin wave theory (equivalent to large-$d$ BOT for $d=\infty$), and large-$N$ SU(4) SB corresponds to this work.
@@
In this work, we develop a systematic large-$N$ expansion based on the $\mathrm{SU}(4)$ coherent state description of a single dimer. This approach generalizes the original dimer problem with two antiferromagnetically coupled $\mathrm{SU}(2)$ spins on each site of the dimer to two interacting $\mathrm{SU}(n)$ spins, with $n\geq 2$. Unlike the traditional SB approximation, which usually employs $\mathrm{SU}(n)$ SBs to represent the $\mathrm{SU}(n)$ spin components [44], we start from $\mathrm{SU}(N=n^{2})$ SBs to include all quantum mechanical states of the single-dimer problem in the manifold of $\mathrm{SU}(N)$ coherent states. This construction defines a family of SU($n$)-invariant models formulated in terms of Schwinger bosons transforming under SU($n^{2}$). Unlike conventional Schwinger boson theories on bipartite lattices [45], where SU($n$) spins (i.e., generators of the SU($n$) group) interact exclusively through antiferromagnetic couplings, our model incorporates additional interaction terms beyond simple antiferromagnetic exchange. To emphasize this distinction, we refer to our approach as the “large-$N$ SU(4) Schwinger boson theory,” explicitly differentiating it from the traditional “large-$N$ SU(2) Schwinger boson theory.” Our theory gives rise to a controlled structure and a $1/N$ expansion of the BOT mean-field approximation, where the boson number constraint is imposed by a Lagrange multiplier.
@@ II $\mathrm{SU}(4)$ coherent-state description of a dimer
where $S$ refers to the spin size, which labels the irrep of $\mathrm{SU}(2)$ that determines the matrix form of the generators given in Eq. (3). Note that $S=1/2$ for the particular case of interest. However, as we mentioned in the Introduction, a large-$N$ approximation based on these SBs is not quantitatively accurate because the path-integral formulation is parametrized in terms of direct products of $\mathrm{SU}(2)$ coherent states. In other words, the intra-dimer entanglement must be built in via quantum fluctuations that favor linear combination of these product states.
@@
provide a faithful representation of generators of $\mathrm{SU}(4)$ in the completely symmetric irreps labelled by the integer $M$ (e.g. $M=1$ for the fundamental irrep of $\mathrm{SU}(4)$). Note that unlike the $\mathrm{SU}(2)$ SBs, these bosons have only a dimer index $j$ because they create all the possible quantum mechanical states of each dimer. In other words, the $\mathrm{SU}(4)$ SBs create coherent states of the $\mathrm{SU}(4)$ Lie algebra (for completely symmetric irreps) that span the CP3 manifold of quantum mechanical states of a four-level quantum mechanical system.
To make contact with our problem of interest, we can regard the single dimer $j$ as the four-level system, set $M=1$ (fundamental irrep of $\mathrm{SU}(4)$) and choose the basis of $\mathrm{SU}(4)$ SBs, such that the $\mu=0$ boson creates the singlet state:
@@
The important difference between the $\mathrm{SU}(2)$ and the $\mathrm{SU}(4)$ SBs is that coherent states of the latter include entangled intra-dimer singlet and triplet states, meaning that the intra-dimer entanglement is already built-in the $\mathrm{SU}(4)$ coherent states of the path-integral formulation and only the inter-dimer entanglement must be generated via the inclusion of quantum fluctuations. As expected from this simple observation, we will see in the next sections that the large-$N$ expansion of the $\mathrm{SU}(4)$ SB theory is indeed much more accurate than the large-$N$ expansion of the $\mathrm{SU}(2)$ SB theory for low-dimensional coupled-dimer systems.
@@ III $\mathrm{SU}(4)$ Schwinger boson theory
In terms of the $\mathrm{SU}(4)$ SB, the intra-dimer Hamiltonian takes the diagonal form:
@@
Given the local constraint in Eq. (5) with $M=1$, the ground state $|\Psi_{0}\rangle$ of ^$\hat{\mathscr{H}}_{0}$ (direct product of singlet states) satisfies $\hat{b}_{j,0}^{\dagger}\hat{b}_{j,0}|\Psi_{0}\rangle=|\Psi_{0}\rangle$ for $\forall j$.
@@
$\hat{\mathscr{H}}^{\prime}$ in Eq. (14) is expressed in a form that is manifestly SU(2) symmetric. Physically, the different terms of $\hat{\mathscr{H}}^{\prime}$ can be interpreted in the following way: $\hat{A}^{\dagger}\hat{S}$ represents the process of destroying two neighboring singlets followed by the creation of two triplets, which together form a singlet (${\rm h.c.}$ corresponds to the inverse process); $\hat{T}^{\dagger}\hat{B}$ the exchange between singlet and triplet neighbours, $:\hat{B}^{\dagger}\hat{B}:$ the exchange between neighboring triplets, and $\hat{A}^{\dagger}\hat{A}$ projects onto singlet formed between two neighboring triplets.
@@ III.1 Large-$N$ generalization
To study this interacting system of SBs, we employ the large-$N$ technique, extending the original $\mathrm{SU}(2)$-invariant model to a broader class of $\mathrm{SU}(n)$ Hamiltonians, expressed in terms SBs with $N=n^{2}$ flavors. Notably, there are multiple ways to construct an $\mathrm{SU}(n)$-invariant interaction term using the generators of the $\mathrm{SU}(N)$ group. Our choice is guided by the following considerations:
  • The generalized Hamiltonian is $SU(n)$-invariant.
  • The form of the generalized Hamiltonian should be preserved across different values of $n$. This ensures a unified approach to analyzing the system for arbitrary $n$.
Regarding the second condition, note that if we were to generalize the model using $\mathrm{SU}(n)$-invariant forms in the $\mathrm{SU}(n)$ spins, the corresponding bond-operator representation would involve a varying number of bond operators depending on the value of $n$, introducing unnecessary complexity in taking the large-$N$ limit.
The Hamiltonian that satisfies both requirements is given by:
where
^$\hat{\mathscr{H}}_{0}=-J\sum_{j}\hat{b}^{\dagger}_{j,0}\hat{b}_{j,0},$ (23)
and
$+$ $\frac{2J^{\prime}}{N}\sum_{j,\delta}\left(\hat{B}_{j,\delta}^{\dagger}\hat{B}_{j,\delta}-\hat{A}_{j,\delta}^{\dagger}\hat{A}_{j,\delta}\right).$
The link operators introduced above are defined as:
$\hat{S}_{j,\delta}$ $=\hat{b}_{j,0}\hat{b}_{j+\delta,0},$ $\hat{T}_{j,\delta}$ $=\hat{b}^{\dagger}_{j,0}\hat{b}_{j+\delta,0},$ (25)
$\hat{A}_{j,\delta}$ $=\sum_{\mu=1}^{N-1}\hat{b}_{j,\mu}\hat{b}_{j+\delta,\mu},$ $\hat{B}_{j,\delta}$ $=\sum_{\mu=1}^{N-1}\hat{b}^{\dagger}_{j,\mu}\hat{b}_{j+\delta,\mu}.$ (26)
Furthermore, we generalize the Schwinger boson number constraint (5) to:
$\sum_{\mu=0}^{N-1}\hat{b}^{\dagger}_{j,\mu}\hat{b}_{j,\mu}=M,$ (27)
where $M$ is an arbitrary integer. For our original problem, we have $N=4$ and $M=1$.
The physical meaning of the generalized model introduced above becomes most transparent for the special case $M=1$. In this scenario, the intra-dimer Hamiltonian, ^$\hat{\mathscr{H}}_{0}$, describes dimers formed by two antiferromagnetically coupled $\mathrm{SU}(n)$ spins:
^$\hat{\mathscr{H}}_{0}\propto\sum_{j}\sum_{\mu=0}^{N-1}\hat{O}_{j,+}^{\mu}\hat{\tilde{O}}_{j,-}^{\mu}+{\rm const.},$ (28)
where $\hat{O}^{\mu}_{j,+}$ represent the generators of the $\mathrm{SU}(n)$ group acting on the “$+$” site of the dimer in the fundamental irreducible representation, and $\hat{\tilde{O}}_{j,-}^{\mu}$ represent those on the “--” site in the conjugate representation. Given the reducible tensor product of these representations, $n\otimes\bar{n}=1\oplus(n^{2}-1)$, the energy spectrum of ^$\hat{\mathscr{H}}_{0}$ consists of a singlet ground state and an $(N-1)$-fold degenerate multiplet transforming as the adjoint irrep of SU($n$), denoted by $\rvert\mu\rangle$. Analogously to the original problem, each energy level can be associated with Schwinger bosons as $\rvert\mu\rangle=\hat{b}_{j,\mu}^{\dagger}\rvert 0\rangle$. In this framework, $\hat{b}_{j,0}$ describes an $\mathrm{SU}(N)$ singlet, while the remaining $N-1$ flavors, $\hat{b}_{j,\mu}$ (with $\mu=1,2,\dots,N-1$), transform according to the self-conjugate adjoint representation of $\mathrm{SU}(N)$. With this identification, the $\mathrm{SU}(n)$ generators are expressed as:
@@
where the indices satisfy $11\leq\mu,\nu,\rho\leq N-$ 1. The coefficients $f_{\mu\nu\rho}$ are the structure constants of the $\mathfrak{su}(n)$ Lie algebra, which are fully antisymmetric under permutation of indices, and $d_{\mu\nu\rho}$ is a completely symmetric third-rank tensor defined through the anticommutator of $\mathrm{SU}(n)$ generators:
$d_{\mu\nu\rho}=2\operatorname{Tr}\left[\{T^{\mu},T^{\nu}\}T^{\rho}\right].$ (31)
In the specific case of $\mathrm{SU}(2)$, $d_{\mu\nu\rho}\equiv 0$ and $f_{\mu\nu\rho}=\epsilon_{\mu\nu\rho}$, thereby recovering the familiar expression for spin operators given in Eq. (12).
Translating the inter-dimer Hamiltonian into the language of ${\mathrm{SU}}(n)$ generators leads to a more intricate structure that inherently includes terms beyond simple inner products of the form $\sum_{\mu=0}^{N-1}\hat{O}_{j,\pm}^{\mu}\hat{\tilde{O}}_{k,\pm}^{\mu}$. Consequently, this generalized family of models departs significantly from conventional ${\mathrm{SU}}(n)$-invariant constructions based on the standard Schwinger boson framework, which relies on the ${\mathrm{SU}}(2)$ representation of spin-$1/2$ operators on bipartite lattices [45]. In fact, there are multiple ways to construct ${\mathrm{SU}}(n)$-invariant models using the $n^{2}-1$ generators of the ${\mathrm{SU}}(n)$ group. The particular approach adopted here introduces a unified representation of link operators that preserves the structure of the inter-dimer Hamiltonian for arbitrary $n$. This construction plays a key role in the broader theoretical framework developed in subsequent sections. Notably, all four link operators—$\hat{S}_{j,\delta}$, $\hat{T}_{j,\delta}$, $\hat{A}_{j,\delta}$, and $\hat{B}_{j,\delta}$—are explicitly ${\mathrm{SU}}(n)$-invariant, thereby ensuring that the generalized Hamiltonian $\hat{\mathscr{H}}$ respects ${\mathrm{SU}}(n)$ symmetry.
For $M>1$, both intra- and inter-dimer Hamiltonians include terms that go beyond simple ${\mathrm{SU}}(n)$-invariant inner products of operators in their higher-dimensional irreps. For example, the energy spectrum of the extended intra-dimer Hamiltonian, ^$\hat{\mathscr{H}}_{0}$, consists of equally spaced energy levels separated by $J$, with a non-degenerate singlet as the ground state. However, this spectrum does not match that of coupled ${\mathrm{SU}}(n)$ generators. The primary motivation for considering large values of $M$ is to enable theoretical approximations, such as semiclassical analyses via the large-$M$ approach (where $M\to\infty$) and the large-$N$ approach discussed herein. The former method captures the classical limit using ${\mathrm{SU}}(N)$ coherent states, with quantum corrections systematically included through a $1/M$ expansion [38, 47, 48]. The latter, the large-$N$ approach, is especially useful for describing quantum states near critical points, such as the QPM-AFM transition explored in this paper. Importantly, $M$ must scale proportionally with $N$, maintaining a finite ratio $\kappa=M/N$, to ensure a nontrivial large-$N$ limit. If this condition is not met, the large-$N$ limit would lead to a vacuum of Schwinger bosons as the ground state, with local creation of Schwinger bosons as the excitations—states that are not of interest here. The distinction between the large-$M$ and large-$N$ approaches is despicted in Fig. 2.